OpenStudy (anonymous):

Solve for this equation algebraically for all values of x and y. y= x^2 + 2x - 8 and y= 2x + 1

OpenStudy (anonymous):

Since the second equation conveniently isolates y for you, we can just go ahead and plug this into your first equation. Therefore: 2x + 1 = x^2 + 2x - 8 Moving all the values to one side gives us: 0 = x^2 - 9 Knowing that (x^2 - y^2) = (x-y)(x+y), this can be broken down into: 0 = (x-3)(x+3) So x = 3 or -3 Now we can just plug these values of x into the second equation and get that y = 7 or -5.

OpenStudy (anonymous):

ohhhh that makes so much sense thank you!